OpenStudy (anonymous):

Factor: (8a^3b^3)-(27c^6d^6)

OpenStudy (anonymous):

\[8a ^{3}b ^{3}-27c ^{6}d ^{6}\]

OpenStudy (aum):

Use the identity: \(\large x^3 - y^3 = (x-y)(x^2 + xy + y^2)\) \(\large (8a^3b^3)-(27c^6d^6) = (2ab)^3 - (3c^2d^2)^3 = ?\)

OpenStudy (anonymous):

yep, follow @aum 's way

OpenStudy (aum):

Put \(\large x = 2ab\) and \(\large y = 3c^2d^2\) on the right hand side of the identity mentioned in my previous reply.